OpenStudy (anonymous):

<1 and <2 are complementary. If m<1 = 63^0 and m<2 = 12 + x, then find the value of x Help Please! :)

ganeshie8 (ganeshie8):

hint : complementary angles add up to `90` degrees

OpenStudy (anonymous):

so it would be like 63^0 +12 +x = 90? If so what do I do for the 0?

ganeshie8 (ganeshie8):

Yes ! \(\large 63^{\circ}\) - \(\large \circ \) is a symbol representing the unit ` degrees`

ganeshie8 (ganeshie8):

\(\large 63^{\circ} + 12+x = 90^{\circ}\)

ganeshie8 (ganeshie8):

solve \(x\)

ganeshie8 (ganeshie8):

you may ignore it when solving the equation - it exists just to tell you that the angle i sin degrees and not in some other crazy units...

ganeshie8 (ganeshie8):

solve it as if the equation is : \(\large 63+ 12+x = 90\)

OpenStudy (anonymous):

so 75x=90??

OpenStudy (anonymous):

wait wait I take it back It would be like x=15

ganeshie8 (ganeshie8):

x=15 is \(\large \color{red}{\checkmark}\)